OpenStudy (brittany):

2x^2+6x-8=72 all i can get up to is 2x2+6x-8-72 = 0

OpenStudy (anonymous):

combine like terms first...so it's 2x^2 + 6x -80

OpenStudy (anonymous):

2x^2 + 6x - 80 = 0 solve the quadratic equation

OpenStudy (anonymous):

now factor out a 2

OpenStudy (anonymous):

2x^2+6x-80=0 2(x+3x-40)=0 2(x+8)(x-5)=0 x can equal -8 or 5. :)

OpenStudy (anonymous):

x^2 + 3x - 50 = 0

OpenStudy (anonymous):

you guys gotta get faster, i'm on a streak. :P

OpenStudy (anonymous):

yea...what some1 said

OpenStudy (anonymous):

yea that made me look bad

OpenStudy (anonymous):

i'm sorry lol. :D

OpenStudy (anonymous):

pellet I gave the wrong answer